Ultraviolet light having a wavelength of 146 nm strikes a metallic surface. Electrons leave the surface with speeds up to 3.15 × 105 m/s. What is the work function, in eV of the metal?
Answer #1

Wavelength of light=146 nm=146*10^-9 m

Speed of light=3*10^8

Plank constant=6.63"10^-34

Frequency of incident light=3*10^8/(146*10^-9)=2*10^15 Hz

Velocity of electron=3.15*10^5 m/s

Mass of electron=9.11*10^-31kg

Kinetic energy of electron =0.5*9.11*10^-31*(3.15*10^5)²=45.2*10^-21 J

Work function=(6.63*10^-34*2*10^15) - (45.2*10^-21)=1326*10^-21-45.2*10^-21=1280.8*10^-21 J=(1280.8*10^-21)/(1.6*10^-19)=8 eV

Therefore work function of the metal is 8eV